OpenStudy (anonymous):

The function f(t) = 3 cos(pi/6t) + 5 represents the tide in Blastic Sea. It has a maximum of 8 feet when time (t) is 0 and a minimum of 2 feet. The sea repeats this cycle every 12 hours. After nine hours, how high is the tide?

OpenStudy (anonymous):

@amistre64 @Nurali @robtobey

OpenStudy (amistre64):

let t=9

OpenStudy (anonymous):

yes i subsituted it into the equation and got 9.701 and thats not a choice

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=+3+cos%289pi%2F6%29+%2B+5 set your calc to radian mode maybe?

OpenStudy (anonymous):

the answer is 5?

OpenStudy (anonymous):

well schitt mines was in degrees thanks

OpenStudy (anonymous):

could you help with another man?

OpenStudy (amistre64):

maybe

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)2 − (sin Θ - cos Θ)2

OpenStudy (amistre64):

its a difference of squares

OpenStudy (anonymous):

Simplify: (sin Θ − cos Θ)^2 − (sin Θ - cos Θ)^2

OpenStudy (amistre64):

x^2 - a^2 = (x-a)(x+a)

OpenStudy (anonymous):

?

OpenStudy (amistre64):

isnt that just zero?

OpenStudy (anonymous):

well thats a choice

OpenStudy (amistre64):

let sin-cos = r r^2 - r^2 = 0

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

is zero the answer i was leaning towards sin^2

OpenStudy (amistre64):

assuming theres no mistype in it

OpenStudy (amistre64):

you have:\[(\color{red}{sin-cos})^2\color{green}-(\color{red}{sin-cos})^2\] \[(\color{red}{r})^2\color{green}-(\color{red}{r})^2\]

OpenStudy (anonymous):

so its basically like adding 1 and negative 1?

OpenStudy (amistre64):

yeh

OpenStudy (anonymous):

k how about another man?

OpenStudy (anonymous):

your a great help

OpenStudy (anonymous):

What is the sine value of 5 pi over 3?

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@amistre64 i need one more plz man

OpenStudy (amistre64):

calculator sin(5pi/3) thats the best i could work it

OpenStudy (amistre64):

its the same as 6pi/3 - pi/3 so id say it the sin of 60 degrees

OpenStudy (anonymous):

i actually figured it out